Aneco an Electrical Company claims that the average lifeof the bulbs it manufactures is 1,200 hours with a standard deviation of 250 hours. If random sample of 100 bulbs is chosen, what is the probability that the sample mean will be.

1
Expert's answer
2022-04-17T17:02:41-0400

"a) P(x>1150) = (x-\mu)/\sigma/\sqrt n = (1150-1200)/250/\sqrt {100} = 0.02275"

"1 - 0.02275 = 0.97725"

97.73%


"b)""P(x<1250) = (x-\mu)/\sigma/\sqrt n = (1250-1200)/250/\sqrt {100} = 0.97725"

"1-0.97725 = 0.02275"

2.28%


c) 0.97725 - 0.02275 = 0.9545

95.45%